As part of an experiment, a chemist measures the temperature of two solutions every minute. At the
beginning of the experiment, the sodium chloride solution’s temperature is 48 degrees higher than
the potassium chloride solution’s temperature. If the sodium chloride solution’s temperature
increases by 1 degree per minute and the potassium chloride’s temperature increases by 7 degrees
per minute, after how many minutes will the two solutions have the same temperature?

Answer by ankor@dixie-net.com(22740)

As part of an experiment, a chemist measures the temperature of two solutions every minute.
At the beginning of the experiment, the sodium chloride solution’s temperature is 48 degrees higher than the potassium chloride solution’s temperature.
If the sodium chloride solution’s temperature increases by 1 degree per minute and the potassium chloride’s temperature increases by 7 degrees per minute, after how many minutes will the two solutions have the same temperature?
;
let m = no. of minutes of the experiment
let t = original temp of the potassium chloride
then
(t+48) = original temp of the sodium chloride
:
7m + t = m + t+48
7m - m = t - t + 48
6m = 48
m = 48/6
m = 8 minutes they will be the same temp
:
:
Check this, find the temp of each after 8 min
Sodium Chloride: 8 + t + 48 = t + 56
potassium Chloride: 7(8) + t = t + 56